OpenStudy (anonymous):

Solve for the roots in the following equation. Hint: Factor both quadratic expressions. (x^4 + 5x^2 - 36)(2x^2 + 9x - 5) = 0

jagr2713 (jagr2713):

2x6+9x5+5x4+45x3−97x2−324x+180

OpenStudy (anonymous):

\[ \left(x^4+5 x^2-36\right) \left(2 x^2+9 x-5\right)=(x-2) (x+2) (x+5) (2 x-1) \left(x^2+9\right)=0 \]